# Scalar Field Dominated Cosmology with Woods-Saxon Like Potential

### v1

@@ Abstract
Dark energy can be characterized by a canonical scalar field, known as quintessence. Quintessence allows for a dynamical equation of state $-1\leq\omega\leq-\frac{1}{3}$. In this work, we consider a quintessence model with a specific potential of the Woods-Saxon form. The model is studied at late phase assuming flat cosmology, and the model parameters are constrained using Type Ia supernova data and Observational Hubble data. In particular we employ Markov Chain Monte Carlo methods for the Bayesian inference of these parameters. We obtain the value of the Hubble constant $H_{0}=68.10\pm 0.74\,kms^{-1}Mpc^{-1}$ and the matter energy density parameter $\Omega_{m_{0}}=0.31\pm 0.01$, which are consistent with the values obtained using base $\Lambda$CDM using the same dataset. Computation of the $\chi^{2}_{min}$, AIC and BIC reveal that this model is preffered according to AIC and $\chi^{2}_{min}$ criteria, while the $\Lambda$CDM is preferred according to BIC. We also show that this model can explain late-time accelerated expansion. Statefinder analysis reveals a Chaplygin gas behaviour in the past, in addition to the quintessence behaviour of the field.
@@ 1 Introduction
Contemporary cosmology attempts to unmask the mystery of dark energy, which drives the (observed) late-time accelerated expansion of the Universe. The initial evidence for this was from the observations of Type Ia Supernovae (SNeIa) as presented in Ref. [1, 2], which confirmed the acceleration of the Universe. According to such observations, the Universe is assumed to be dominated by a mysterious component called dark energy, constituting about $70\%$ of the present Universe, the rest consisting of matter, of which normal baryonic matter is only around $4\%$, and the remaining is in the form of ”invisible” dark matter. This inference has also been confirmed by observations of the CMB [3, 4] and Baryonic Acoustic Osillations [5, 6].
@@
Dark energy is characterized by a perfect fluid with a negative equation of state, $\omega=P/\rho$. One of the simplest models for dark energy is the cosmological constant $\Lambda$. The cosmological constant has a characteristic equation of state $\omega=-1$. This is the currently accepted model of dark energy, by the standard $\Lambda$CDM. Physical explanation of the cosmological constant may be attributed to the zero-point quantum vacuum fluctuations in the Universe [7], given by $\Lambda=8\pi G\rho_{vac}/c^{2}$. However, theoretical predictions of $\rho_{vac}$ do not match with the current observations, by several orders of magnitude [8, 9]. This cosmological constant problem, suggests that the mentioned explanation for the cosmological constant is unsatisfactory. In addition to this, $\Lambda$CDM encounters problems such as the Cosmological Coincidence Problem [10] and the tensions in the inferred values of various parameters between low redshift data [11, 12, 13] and high redshift data [14, 15, 16]. Attempts to resolve these issues have been ongoing and many possible solutions have been suggested in the form of alternate cosmological models. These include the modified gravity models (Ref. [17, 18, 19, 20, 21]), which modify the curvature term in Einstein Field Equations, and modified (or dynamical) dark energy models which modify the form of the energy-momentum tensor in the field equations [22, 23, 24, 25, 26, 27, 28, 29, 30, 31].
@@
Such kinds of modified dark energy models characterize dark energy as a specific form of matter, such as quintessence[32], k-essence [33] or the Chaplygin gas [34]. The quintessence is a particular model in which a scalar field is slow rolling down a potential that is similar to that of early Universe inflation models [35]. Reviews of various quintessence models can be found in Ref. [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. Another kind of scalar field model is k-essence, which, unlike quintessence, does not have a canonical kinetic term. These models have also been studied extensively [48, 49, 50, 51, 52] much like quintessence. The advantage of these scalar field models is that some of these models can address the fine tuning and coincidence problem, all the while explaining the late-time cosmic acceleration. These models are distinguished by their dynamical equation of state $\omega$, which is not the case for $\Lambda$CDM.
@@
The aim of this paper is to construct a quintessence model with a potential which has a similar form as that of the Woods-Saxon potential of nuclear physics. In Sec. 2, we present our proposed model, which will be referred to as the Woods-Saxon Quintessence or WSQ hereafter. It is essential to emphasize that while we draw inspiration from the form of the Woods-Saxon potential, this model is entirely distinct in its interpretation within the context of quintessence cosmology. In Sec. 3, we employ Bayesian inference to constrain the model parameters using observational data, specifically the Type Ia supernova data and observational Hubble parameter data. In Sec. 4, we investigate the evolution of various cosmological parameters within the framework of the WSQ model. Finally, we conclude our work in Sec. 5.
@@ 2 Woods-Saxon Quintessence
$S=\int d^{4}x\sqrt{-g}\left[\frac{1}{2}\frac{R}{\kappa^{2}}-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-V(\phi)\right]+S_{m},$ (1)
@@
$\omega_{\phi}=\frac{P_{\phi}}{\rho_{\phi}}=\frac{\frac{\dot{\phi}^{2}}{2}-V(\phi)}{\frac{\dot{\phi}^{2}}{2}+V(\phi)}.$ (2)
@@
$\dot{\rho_{\phi}}+3H(\rho_{\phi}+P_{\phi})=0,$ (3)
$\dot{\rho_{m}}+3H\rho_{m}=0,$ (4)
where the dark matter is assumed to be pressureless, giving an equation of state $\omega_{m}=0$. Introducing the forms of $\rho_{\phi}$ and $P_{\phi}$ in Eq. (3), we obtain the Klein-Gordon equation,
$\ddot{\phi}+3H\dot{\phi}+V_{,\phi}=0,$ (5)
@@
$3H^{2}=\kappa^{2}\left[\frac{\dot{\phi}^{2}}{2}+V(\phi)+\rho_{m}\right],$ (6)
$2\dot{H}=-\kappa^{2}[\dot{\phi}^{2}+\rho_{m}].$ (7)
@@
$V(\phi)=\frac{2V_{0}}{e^{\kappa(\phi-\phi_{0})}+1},$ (8)
where $V_{0}$ and $\phi_{0}$ are the values of the potential energy and scalar field at present time, i.e., at redshift $z=0$, respectively. Analytical solution to Eq. (7) is challenging with the potential $V(\phi)$. In order to solve Eq. (7) numerically, we introduce three dimensionless parameters,
$x=\frac{\dot{\phi}\kappa}{\sqrt{6}H_{0}},\hskip 14.45377pty=\frac{\sqrt{V(\phi)}\kappa}{\sqrt{3}H_{0}},\hskip 14.45377pth=\frac{H}{H_{0}},$ (9)
where $H_{0}$ is the Hubble parameter at present. Taking the derivatives of $x$, $y$ and $h$ with $z$ and using Eq. (5), (6) and (7), we obtain three coupled differential equations,
$\frac{dx}{dz}=\frac{3x}{1+z}-\sqrt{\frac{3}{2}}\frac{y^{2}}{h(1+z)}\left[1-\frac{1}{2}\left(\frac{y}{y_{0}}\right)^{2}\right],$ (10)
$\frac{dy}{dz}=\sqrt{\frac{3}{2}}\frac{xy}{h(1+z)}\left[1-\frac{1}{2}\left(\frac{y}{y_{0}}\right)^{2}\right],$ (11)
$\frac{dh}{dz}=\frac{3}{2}\frac{h^{2}+x^{2}-y^{2}}{h(1+z)}.$ (12)
@@
$h^{2}=x^{2}+y^{2}+\Omega_{m},$ (13)
where we have defined $\Omega_{m}=\rho_{m}/\rho_{c_{0}}=\rho_{m}\kappa^{2}/3H_{0}^{2}$, where $\rho_{c_{0}}=3H_{0}^{2}/\kappa^{2}$ is the present value of the critical density. We define the energy density parameter of the field , $\Omega_{\phi}$ using Eq. (9) in terms of the variables $x$ and $y$ as,
$\Omega_{\phi}=\frac{\kappa^{2}\rho_{\phi}}{3H_{0}^{2}}=x^{2}+y^{2}.$ (14)
The equation of state can be written using Eq. (2) as,
$\omega_{\phi}=\frac{x^{2}-y^{2}}{x^{2}+y^{2}}.$ (15)
@@ 3 Data Analysis
In order to perform the data analysis, we have to solve the coupled differential equations presented in Eq. (10)-(12), which require the initial conditions $x_{0}$, $y_{0}$ and $h_{0}$. These parameters are constrained by the equation $1=x_{0}^{2}+y_{0}^{2}+\Omega_{m_{0}}$, where $h_{0}=1$. Thus, the WSQ model has effectively three independent free parameters $H_{0}$, $\Omega_{m_{0}}$ and $y_{0}$. In order to constrain these parameters, we employ Markov Chain Monte Carlo methods for Bayesian inference [53]. The primary input for the Bayesian inference is the prior range for each model parameter. We assume a sufficiently large prior range for $H_{0}$ between $40$ and $100$, and 00 to $1$ for $\Omega_{m_{0}}$. The prior range for $y_{0}$ is assumed to be $-1$ to $1$ as the present density parameters of the components add to unity. The parameters are constrained using Observational Hubble Data (OHD) and Type Ia Supernova (SNIa) data. The datasets we used are that from [54] for OHD and the Pantheon+ sample [55] for SNIa.
Figure 1: The 2D confidence contours for 68% and 95% probabilities and 1D posterior distribution of the model parameters using SNIa+OHD datasets for $\Lambda$CDM and the WSQ model. The chosen prior ranges are $40\leq H_{0}\leq 100$, $0\leq\Omega_{m_{0}}\leq 1$, $-20\leq M\leq-18$ and $-1\leq y_{0}\leq 1$
The OHD dataset consists of $77$ redshift vs Hubble parameter values in the redshift range $0\leq z\leq 2.36$, obtained from cosmic chronometers, BAO signal in galaxy distribution and BAO signals in the $Ly\alpha$ forest distribution. The Pantheon+ data comprises the sample of $1701$ Type Ia supernova light curves of 1550 distinct supernovae discovered in the redshift range $0\leq z\leq 2.3$ [56]. The apparent magnitude of SNIa, $m$ is obtained from the expression,
$m(z)=5\,log_{10}\left[\frac{d_{L}(z)}{Mpc}\right]+M+25,$ (16)
@@
$d_{L}(z)=c(1+z)\int_{0}^{z}\frac{dz^{\prime}}{H(z^{\prime})},$ (17)
@@
$\chi^{2}_{OHD}=\sum_{i=1}^{77}\left[\frac{H_{obs}(z_{i})-H_{th}(z_{i},H_{0},\Omega_{m_{0}},y_{0})}{\sigma(z_{i})}\right]^{2},$ (18)
@@
$\chi^{2}_{SNIa}=\Delta\vec{D}^{T}C^{-1}_{stat+syst}\Delta\vec{D},$ (19)
@@
$\chi^{2}_{total}=\chi^{2}_{OHD}+\chi^{2}_{SNIa}.$ (21)
The parameter set ($H_{0}$, $\Omega_{m_{0}}$, $M$, $y_{0}$) that minimises the $\chi^{2}_{total}=-2\,log_{e}(\mathcal{L})$, where $\mathcal{L}$ is the Gaussian likelihood, is the best-fit parameters. We utilize the information criteria such as Akaike Information Criterion (AIC) and Bayesian Information Criterion (BIC) for the model comparison. We use the definitions of AIC and BIC as,
$AIC=n\,log_{e}\left(\frac{\chi^{2}}{n}\right)+2k,$ (22)
$BIC=n\,log_{e}\left(\frac{\chi^{2}}{n}\right)+k\,log_{e}(n).$ (23)
We used a well tested and open source PYTHON implementation of the affine-invariant ensemble sampler for MCMC proposed by Goodman and Weare called emcee for parameter inference. More details about emcee and its implementation can be found in Ref. [57, 58]. We also employed the multiprocessing module from PYTHON standard library and the numba module to reduce the computation time. The best-fit values of the model parameters are presented in Table 1 and the marginal likelihood of the model parameters are presented in Fig. 1. The minimum $\chi^{2}$ obtained for the WSQ model with the SNIa + OHD dataset is $\chi^{2}_{min}=1813.94$ while for $\Lambda$CDM, we obtained $\chi^{2}_{min}=1818.39$. This suggests that the WSQ model is preffered over the $\Lambda$CDM according to the $\chi^{2}_{min}$ criteria.
Data SNIa + OHD
$H_{0}$ $68.1041\pm 0.7374$
$\Omega_{m_{0}}$ $0.3059\pm 0.0129$
$M$ $-19.4126\pm 0.0212$
$y_{0}$ $0.8251\pm 0.0072$
@@
A similar trend is followed with the AIC value, with $\Delta$AIC $=2.36$. However, BIC tends to penalize the extra parameter to a greater extent, so the $\Lambda$CDM is preferred over the WSQ model according to the $\Delta$BIC criteria.
Model $\chi^{2}_{min}$ AIC BIC
$\Lambda$CDM 1818.3949 45.9428 62.3926
WSQ 1813.9435 43.5849 65.5179
Table 2: The $\chi^{2}_{min}$, Akaike Information Criterion (AIC) and Bayesian Information Criterion (BIC) obtained for the two models for the SNIa + OHD dataset.
The estimated value for the Hubble constant $H_{0}=68.10\pm 0.74$, which is in close agreement with the Planck 2018 results [15]. The current value of the dark energy density parameter is computed as $\Omega_{\phi_{0}}=0.6941\pm 0.0129$. With the estimated parameters, we analyse the background evolution of cosmological parameters within the WSQ background.

### v2

@@ Abstract
Dark energy can be characterized by a canonical scalar field, known as quintessence. Quintessence allows for a dynamical equation of state $-1\leq\omega\leq-\frac{1}{3}$. A previous study by Oikonomou and Chatzarakis have shown that a scalar field model with a Woods-Saxon like potential can successfully explain the early inflation. In this work, we consider a quintessence model with a potential of similar form to explain the late time acceleration. The model is studied at late phase assuming flat cosmology, and the model parameters are constrained using Type Ia supernova data and Observational Hubble data. In particular we employ Markov Chain Monte Carlo methods for the Bayesian inference of these parameters. We obtain the value of the Hubble constant $H_{0}\sim 68\text{ km s}^{-1}\text{Mpc}^{-1}$ and the matter energy density parameter $\Omega_{m_{0}}\sim 0.30$, which are in close agreement with the values obtained from the Planck CMB data, assuming the $\Lambda$CDM model. Computation of the $\chi^{2}_{min}$, AIC and BIC reveal that this model is slightly preferred according to AIC and $\chi^{2}_{min}$ criteria, while the $\Lambda$CDM is preferred according to BIC. We demonstrate that the model possesses a stable attractor in the asymptotic future, which confirms the dynamical stability of the model. Thus, this model may be considered as a potential alternative to the $\Lambda$CDM.
@@ 1 Introduction
Contemporary cosmology attempts to unmask the mystery of dark energy, which drives the (observed) late-time accelerated expansion of the Universe. The initial evidence for this was from the observations of Type Ia Supernovae (SNeIa) as presented in Ref. [1, 2], which confirmed the acceleration of the Universe. According to such observations, the Universe is assumed to be dominated by a mysterious component called dark energy, constituting about $70\%$ of the present Universe, the rest consisting of matter, of which normal baryonic matter is only around $4\%$, and the remaining is in the form of ”invisible” dark matter. This inference has also been confirmed by other observations such that of the CMB [3, 4, 5, 6] and Baryonic Acoustic Osillations [7, 8, 9].
@@
Dark energy is characterized by a perfect fluid with a negative equation of state, $\omega=P/\rho$. One of the simplest models for dark energy is the cosmological constant $\Lambda$[10, 11]. The cosmological constant has a characteristic equation of state $\omega=-1$. This is the currently accepted model of dark energy, by the standard $\Lambda$CDM. Physical explanation of the cosmological constant may be attributed to the zero-point quantum vacuum fluctuations in the Universe [12], given by $\Lambda=8\pi G\rho_{vac}/c^{2}$. However, theoretical predictions of $\rho_{vac}$ do not match with the current observations, by several orders of magnitude [13, 14]. This cosmological constant problem, suggests that the mentioned explanation for the cosmological constant is unsatisfactory. In addition to this, $\Lambda$CDM encounters problems such as the Cosmological Coincidence Problem [15] and the tensions in the inferred values of various parameters between low redshift data [16, 17, 18] and high redshift data [19, 20]. Attempts to resolve these issues have been ongoing and many possible solutions have been suggested in the form of alternate cosmological models. These include the modified gravity models (Ref. [21, 22, 23, 24, 25]), which modify the curvature term in Einstein Field Equations, and modified (or dynamical) dark energy models which modify the form of the energy-momentum tensor in the field equations [26, 27, 28, 29, 30, 31, 32, 33, 34, 35].
@@
Such kinds of modified dark energy models characterize dark energy as a specific form of matter, such as quintessence [36], k-essence [37] or the Chaplygin gas [38]. The quintessence is a particular model in which a scalar field is slow rolling down a potential that is similar to that of early Universe inflation models [39]. Reviews of various quintessence models can be found in Ref. [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. Another kind of scalar field model is k-essence, which, unlike quintessence, does not have a canonical kinetic term. These models have also been studied extensively [52, 53, 54, 55] much like quintessence. The advantage of these scalar field models is that some of these models can address the fine tuning and coincidence problem, all the while explaining the late-time cosmic acceleration. These models are distinguished by their dynamical equation of state $\omega$, which is not the case for $\Lambda$CDM.
@@
Quintessence models, as mentioned, have been explored as alternatives to address the shortcomings of the standard $\Lambda$CDM. Various quintessence models have been studied in literature, giving specific forms to the potential energy of the scalar field[56, 57, 58, 59, 60, 61, 62, 63]. But it is important that the model exhibits a stable attractor solution in the asymptotic future, so that a large assortment of initial conditions can lead to the same stable future. Else it would have the same fine tuning problem as the concordance model. In addition, the considered potential must be able to reproduce the cosmic acceleration. Dynamical system analysis is necessary to verify the existence of such a stable attractor. This technique is widely used to cosmology to analyse the stability of a model[64], especially for scalar field models[65]. The stability of critical (or ”fixed”) points are examined mainly using the linear stability theorem,[66] and to analyse non-hyperbolic critical points, centre manifold theorem[67, 68] is employed.
The aim of this paper is to construct a quintessence model with a potential which has a similar form as that of the Woods-Saxon potential of nuclear physics. It is essential to emphasize that while we draw inspiration from the form of the Woods-Saxon potential, this model is entirely distinct in its interpretation within the context of quintessence cosmology. A similar scalar field approach with the Woods-Saxon potential has shown promising results as an inflationary model for the primordial era of the Universe. The authors of [69] were able to successfully realize an inflationary phase along with a consistent reheating epoch with the scalar field model, which is compatible with the Planck CMB data. This study aims to test the phenomenological implications in the late-time accelerated expansion of the Universe with a Woods-Saxon like potential in a quintessence scenario. We constrain the model with observations, and study the evolution of the Universe in this model. In our dynamical system analysis, we obtain one non-hyperbolic point which had to be studied using the centre manifold approach. This approach makes it possible to study the dynamics of the system near the critical point by reducing its dimensionality. The analysis reveals the existence of a stable attractor solution, indicating that the Universe is dynamically stable within the framework of this model.
This paper is organized as follows: In Sec. 2, we present our proposed model, which will be referred to as the Woods-Saxon Quintessence or WSQ hereafter. In Sec. 3, we employ Bayesian inference to constrain the model parameters using observational data, specifically the Type Ia supernova data and observational Hubble parameter data. In Sec. 4, we investigate the evolution of various cosmological parameters within the framework of the WSQ model. In Sec. 5, we test the dynamical stability of the proposed WSQ model by performing dynamical system analysis. Finally, we conclude our work in Sec. 6.
@@ 2 Woods-Saxon Quintessence
$S=\int d^{4}x\sqrt{-g}\left[\frac{1}{2}\frac{R}{\kappa^{2}}-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-V(\phi)\right]+S_{m},$ (2.1)
@@
$\omega_{\phi}=\frac{P_{\phi}}{\rho_{\phi}}=\frac{\frac{\dot{\phi}^{2}}{2}-V(\phi)}{\frac{\dot{\phi}^{2}}{2}+V(\phi)}.$ (2.2)
@@
$\dot{\rho_{\phi}}+3H(\rho_{\phi}+P_{\phi})=0,$ (2.3)
$\dot{\rho_{m}}+3H\rho_{m}=0,$ (2.4)
where the dark matter is assumed to be pressureless, giving an equation of state $\omega_{m}=0$. Introducing the forms of $\rho_{\phi}$ and $P_{\phi}$ in Eq. (2.3), we obtain the Klein-Gordon equation,
$\ddot{\phi}+3H\dot{\phi}+V_{,\phi}=0,$ (2.5)
@@
$3H^{2}=\kappa^{2}\left[\frac{\dot{\phi}^{2}}{2}+V(\phi)+\rho_{m}\right],$ (2.6)
$2\dot{H}=-\kappa^{2}[\dot{\phi}^{2}+\rho_{m}].$ (2.7)
@@
$V(\phi)=\frac{2V_{0}}{e^{\kappa(\phi-\phi_{0})}+1},$ (2.8)
where $V_{0}$ and $\phi_{0}$ are the values of the potential energy and scalar field at present time, i.e., at redshift $z=0$, respectively. Analytical solution to Eq. (2.7) is challenging with the potential $V(\phi)$. In order to solve Eq. (2.7) numerically, we introduce three dimensionless parameters,
$x=\frac{\dot{\phi}\kappa}{\sqrt{6}H_{0}},\hskip 14.45377pty=\frac{\sqrt{V(\phi)}\kappa}{\sqrt{3}H_{0}},\hskip 14.45377pth=\frac{H}{H_{0}},$ (2.9)
where $H_{0}$ is the Hubble parameter at present. Taking the derivatives of $x$, $y$ and $h$ with $z$ and using Eq. (2.5), (2.6) and (2.7), we obtain three coupled differential equations,
$\frac{dx}{dz}=\frac{3x}{1+z}-\sqrt{\frac{3}{2}}\frac{y^{2}}{h(1+z)}\left[1-\frac{1}{2}\left(\frac{y}{y_{0}}\right)^{2}\right],$ (2.10)
$\frac{dy}{dz}=\sqrt{\frac{3}{2}}\frac{xy}{h(1+z)}\left[1-\frac{1}{2}\left(\frac{y}{y_{0}}\right)^{2}\right],$ (2.11)
$\frac{dh}{dz}=\frac{3}{2}\frac{h^{2}+x^{2}-y^{2}}{h(1+z)}.$ (2.12)
@@
$h^{2}=x^{2}+y^{2}+\Omega_{m},$ (2.13)
where we have defined $\Omega_{m}=\rho_{m}/\rho_{c_{0}}=\rho_{m}\kappa^{2}/3H_{0}^{2}$, where $\rho_{c_{0}}=3H_{0}^{2}/\kappa^{2}$ is the present value of the critical density. We define the energy density parameter of the field , $\Omega_{\phi}$ using Eq. (2.9) in terms of the variables $x$ and $y$ as,
$\Omega_{\phi}=\frac{\kappa^{2}\rho_{\phi}}{3H_{0}^{2}}=x^{2}+y^{2}.$ (2.14)
The equation of state can be written using Eq. (2.2) as,
$\omega_{\phi}=\frac{x^{2}-y^{2}}{x^{2}+y^{2}}.$ (2.15)
@@ 3 Data Analysis
In order to perform the data analysis, we have to solve the coupled differential equations presented in Eq. (2.10)-(2.12), which require the initial conditions $x_{0}$, $y_{0}$ and $h_{0}$. These parameters are constrained by the equation $1=x_{0}^{2}+y_{0}^{2}+\Omega_{m_{0}}$, where $h_{0}=1$. Thus, the WSQ model has effectively three independent free parameters $H_{0}$, $\Omega_{m_{0}}$ and $y_{0}$. In order to constrain these parameters, we employ Markov Chain Monte Carlo methods for Bayesian inference [70]. The primary input for the Bayesian inference is the prior range for each model parameter. We assume a sufficiently large prior range for $H_{0}$ between $40$ and $100$, and 00 to $1$ for $\Omega_{m_{0}}$. The prior range for $y_{0}$ is assumed to be 00 to $1$ as the present density parameters of the components add to unity, and $y$ must be positive by definition. The parameters are constrained using Observational Hubble Data (OHD) and Type Ia Supernova (SNIa) data. The dataset we used for the OHD data is from Ref. [71]. We considered two different datasets for the SNIa data - The Pantheon+ dataset[72] and the five year Dark Energy Survey (DES) data [73, 74]. The OHD dataset consists of $77$ redshift vs Hubble parameter values in the redshift range $0\leq z\leq 2.36$, obtained from cosmic chronometers, BAO signal in galaxy distribution and BAO signals in the $Ly\alpha$ forest distribution. The Pantheon+ data comprises the sample of $1701$ Type Ia supernova light curves of 1550 distinct supernovae discovered in the redshift range $0\leq z\leq 2.3$ [75]. We will refer to the combination of Pantheon+ and OHD dataset as Pantheon+OHD. The DES data consists of a total sample of $1829$ comprising of $1635$ DES SNe and $194$ external low-redshift samples in the redshift range $0.025\leq z\leq 1.13$. We will refer to the combination of DES and OHD datasets as the DES+OHD. The SNIa datasets consist of redshift $z$ against apparent magnitude ($m$) obtained from their respective surveys. The apparent magnitude of SNIa, $m$ is obtained from the expression,
$m(z)=5\,log_{10}\left[\frac{d_{L}(z)}{Mpc}\right]+M+25,$ (3.1)
@@
$d_{L}(z)=c(1+z)\int_{0}^{z}\frac{dz^{\prime}}{H(z^{\prime})},$ (3.2)
@@
$\chi^{2}_{OHD}=\sum_{i=1}^{77}\left[\frac{H_{obs}(z_{i})-H_{th}(z_{i},H_{0},\Omega_{m_{0}},y_{0})}{\sigma(z_{i})}\right]^{2},$ (3.3)
@@
$\chi^{2}_{SNIa}=\Delta\vec{D}^{T}C^{-1}_{stat+syst}\Delta\vec{D},$ (3.4)
@@
$\chi^{2}_{total}=\chi^{2}_{OHD}+\chi^{2}_{SNIa}.$ (3.6)
Figure 1: The 2D confidence contours for 68% and 95% probabilities and 1D posterior distribution of the model parameters using SNIa+OHD datasets for $\Lambda$CDM and the WSQ model. The chosen prior ranges are $40\leq H_{0}\leq 100$, $0\leq\Omega_{m_{0}}\leq 1$, $-20\leq M\leq-18$ and $0\leq y_{0}\leq 1$.
The parameter set ($H_{0}$, $\Omega_{m_{0}}$, $M$, $y_{0}$) that minimises the $\chi^{2}_{total}=-2\,\text{ln}(\mathcal{L})$, where $\mathcal{L}$ is the Gaussian likelihood, is the best-fit parameters. We utilize the information criteria such as Akaike Information Criterion (AIC) and Bayesian Information Criterion (BIC) for the model comparison. We use the definitions of AIC and BIC as,
$AIC=n\,\text{ln}\left(\frac{\chi^{2}}{n}\right)+2k,$ (3.7)
$BIC=n\,\text{ln}\left(\frac{\chi^{2}}{n}\right)+k\,\text{ln}(n).$ (3.8)
We used a well tested and open source PYTHON implementation of the affine-invariant ensemble sampler for MCMC proposed by Goodman and Weare called emcee for parameter inference. More details about emcee and its implementation can be found in Ref. [76, 77]. We also employed the multiprocessing module from PYTHON standard library and the numba module to reduce the computation time. The best-fit values of the model parameters are presented in Table 1 and the marginal likelihood of the model parameters are presented in Fig. 1. The minimum $\chi^{2}$ obtained for the WSQ model with the SNIa + OHD dataset is lower than that obtained for $\Lambda$CDM. This suggests that the WSQ model is preffered over the $\Lambda$CDM according to the $\chi^{2}_{min}$ criteria.
Data Pantheon + OHD DES + OHD
$H_{0}$ $68.1041\pm 0.7374$ $68.2021\pm 0.7374$
$\Omega_{m_{0}}$ $0.3059\pm 0.0129$ $0.3040\pm 0.0124$
$M$ $-19.4126\pm 0.0212$ $-19.3789\pm 0.0194$
$y_{0}$ $0.8251\pm 0.0072$ $0.8262\pm 0.0069$
@@
A similar trend is followed with the Akaike Information Criterion (AIC) value, however, the Bayesian Information Criterion (BIC) tends to penalize the extra parameter to a greater extent, so the $\Lambda$CDM is preferred over the WSQ model according to the $\Delta$BIC criteria. The information criteria we used to compare the fit of the models is given in Table 2.
Dataset Model $\Delta\chi^{2}_{min}$ $\Delta$AIC $\Delta$BIC
Pantheon+OHD $\Lambda$CDM 0 0 0
WSQ 4.4514 2.3579 -3.1253
DES+OHD $\Lambda$CDM 0 0 0
WSQ 3.6300 2.0642 -3.4885
Table 2: The $\Delta\chi^{2}_{min}$, $\Delta AIC$ and $\Delta BIC$ obtained for the two models for the SNIa + OHD datasets.
The estimated value for the Hubble constant, from both the datasets, is in close agreement with the Planck 2018 results [20]. The current value of the dark energy density parameter is computed as $\Omega_{\phi_{0}}=0.6941\pm 0.0129$ with the Pantheon+OHD dataset and $\Omega_{\phi_{0}}=0.6960\pm 0.0124$ with the best fit values obtained for the DES+OHD dataset. With the estimated parameters, we analyse the background evolution of cosmological parameters within the WSQ background.